Give examples of 3 X 3 matrices A,B, and C satisfying the following conditions: A^2 = 2A, B^2 = 2B, C^2 =/= 2C, det(A) = 0, det(B) = det(C) = 8.
1
Expert's answer
2020-02-14T12:56:32-0500

"A=\begin{bmatrix}
 0 & 0 & 0 \\
 0 & 0 & 0 \\
 0 & 0 & 0
\end{bmatrix}"


"B=\begin{bmatrix}
 2 & 0 & 0 \\
 0 & 2 & 0 \\
 0 & 0 & 2
\end{bmatrix}"


"C=\begin{bmatrix}
 0 & 2 & 0 \\
 0 & 0 & 2 \\
 2 & 0 & 0
\end{bmatrix}"